OpenStudy (anonymous):

Find the constant of variation for the quadratic variation. 8 16 24 40

OpenStudy (anonymous):

x = 2 / 3 / 4 / 5 / 6 y = 32/72/128/200/288

OpenStudy (anonymous):

r u almost done?

OpenStudy (mathmale):

Howdy, Cookie! We use different models for different types of variation. For example, linear variation is associated with the formula y=ax, or the more familiar y=mx+b (the equation of a straight line). Cubic variation: y=a*x^3. In the present case we're discussing quadratic variation; perhaps that will ring a bell with you, reminding you that y=ax^2+bx+c is the general quadratic function. Now in y our math problem, we're told that this is a case of quadratic variation. Use the model y=a*x^2. For example, we know that if x=2, y =32. Mind substituting those two values into y=a*x^2 and solving for y? Then you could re-write y=a*x^2 substituting this value for a. Then check thisd value by substituting x=3, y=72, and see whether the resulting equation is true or not. If it is, your a value is correct.

OpenStudy (mathmale):

So: Please bring out pencil and paper and find a.

OpenStudy (anonymous):

i just wanna check my asnwers

OpenStudy (mathmale):

There's only 1 "answer" to this problem, only one a value. Could y ou find it, please?

OpenStudy (anonymous):

these are the choices

OpenStudy (anonymous):

8,16 ,24 ,40

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Cookie, I'm a math teacher, not a student. I'd be very interested in helping you to understand the concept behind this math question. If you just want an answer, I'm the wrong person to ask. Mind doing the work I've suggested and sharing it with me?

OpenStudy (anonymous):

i got 16